OpenStudy (anonymous):

Which of the following is a factor of x2- 21x + 110? A) x + 11 B) x - 10 C) x - 5 D) None of the above

OpenStudy (anonymous):

x-10

OpenStudy (anonymous):

see how..

OpenStudy (anonymous):

Which is a factor of 2x2+ 8x - 90 ? A) x + 2 B) x - 15 C) x + 9 D) x - 4

OpenStudy (anonymous):

x2-21x+110=x2-11x-10x+110 => =x(x-11)-10(x-11) => =(x-11)(x-10) so x-10 is a fector..